OpenStudy (anonymous):

solve 3+sqrt z-11=sqrt z+4

OpenStudy (anonymous):

post the question clearly using the equation button below | | v

OpenStudy (anonymous):



OpenStudy (anonymous):

alright. square both sides and post what you get.

OpenStudy (anonymous):

idk how to use equation editor lol i tried

OpenStudy (anonymous):

i dont know how to square both sides... i dont even know what that means

OpenStudy (anonymous):

z-22 and z+8 ??

OpenStudy (anonymous):

no no. square simply means taking a number and multiplying it by itself. for example, 4 multiplied by itself gives 16. that is, 4 x 4 = 16 therefore 16 is the square of 4. conversely, 4 is the square root of 16 since 4 is the root or the starting point.

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

so its-120 and 16

OpenStudy (anonymous):

i need to get my calculator out lol

OpenStudy (anonymous):

so what fdo I do with the 3 ?

OpenStudy (anonymous):

so when you have \[(3\sqrt{z-11})^2 = 3^2 \times (\sqrt{z-11})^2\]

OpenStudy (anonymous):

now, remember that \[4 = \sqrt {16}\] \[(\sqrt{16})^2 = 4^2 = 16\]

OpenStudy (anonymous):

\[3 + \sqrt{z+11} = \sqrt{z+4}\] is this question?

OpenStudy (anonymous):

yes sheg

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait, is 3 added or multiplied?

OpenStudy (anonymous):

oh pellet its added my bad

OpenStudy (anonymous):

i didnt realize i typed it wrong until just now Im so sorry

OpenStudy (anonymous):

i like how they bleep out my pellet with pellet fluttering retarded

OpenStudy (anonymous):

lmfao ha ha

OpenStudy (anonymous):

\[9 + 6\sqrt{z-11} + z-11 = z+4\] \[6\sqrt{z-11} = 6\] \[\sqrt{z-11} = 1\] \[{z-11} = 1\] \[{z} = 12\]

OpenStudy (anonymous):

well where did the 3 go?

OpenStudy (anonymous):

so u have z=12

OpenStudy (anonymous):

awesome i only have like 5 more ?'s

OpenStudy (anonymous):

(a+b)^2= a^2+(2*a*b)+B^2 this is how to solve a polynomial equation of a second degree so in your equation is a=3 and b=root squirt (z-11) you apply the formula and you will know where did the 3 go